OpenStudy (sparklyme):

I'm having some trouble with this problem. I know the formulas, but I'm not sure how to insert them into this equation. Can I please have some help? I am looking for an explanation, not just an answer. A boy and the skateboard on which he is standing have a combined mass of 60 kg. He throws a 1kg rock forward at 20 m/sec. At what speed will the boy roll backward? I know that Speed = Distance/Time but I'm kind of having trouble because mass is in the problem but not in the equation.

OpenStudy (osprey):

This question seems to be about CONSERVATION OF LINEAR MOMENTUM. The TOTAL momentum BEFORE throwing is zero, 0. (Nothing is moving) The boy's mass is 59kg. The momentum of the ROCK AFTER throwing is 1 x 20 kgm/s = 20kgm/s So, the BOY'S momentum AFTER throwing must be equal and opposite (that's why it's a recoil). 59 times speed = 20 speed = 10/59 = 0.33 m/s ish .. I HOPE http://perendis.webs.com

OpenStudy (osprey):

This question seems to be about CONSERVATION OF LINEAR MOMENTUM. The TOTAL momentum BEFORE throwing is zero, 0. (Nothing is moving) The boy's mass is 59kg. The momentum of the ROCK AFTER throwing is 1 x 20 kgm/s = 20kgm/s So, the BOY'S momentum AFTER throwing must be equal and opposite (that's why it's a recoil). 59 times speed = 20 speed = 20/59 = 0.33 m/s ish .. I HOPE http://perendis.webs.com

OpenStudy (osprey):

second post supercedes first post. error therein. corrected hopefully.

OpenStudy (sapphiremoon):

psst... @osprey... you can delete your old replies... button bottom right corner... ;)

OpenStudy (blazeryder):

If you are still having trouble understanding check this link out: http://aplusphysics.com/courses/honors/momentum/conservation.html

OpenStudy (osprey):

thanks sapphire ... more just Site weariness than anything else ... twinkle on

OpenStudy (kunaliiserb):